OpenStudy (anonymous):

Evaluate the integral of sinx(1+cos³x) dx from (0, π)

OpenStudy (anonymous):

let u= cos(x)

OpenStudy (anonymous):

du= -sin(x) dx sin(x) dx= -du

OpenStudy (anonymous):

Change limits when x=pi, u= -1 when x=0, u=1

OpenStudy (anonymous):

\[= \int\limits_{1}^{-1} -(1+u^3) du = \int\limits_{-1}^{1} (1+u^3) du \]

OpenStudy (anonymous):

you can do the rest, its easy.

OpenStudy (anonymous):

=2

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

wait is the integral of 1 = u?